OpenStudy (shamil98):

\[\int\limits_{}^{} \frac{ x^2 + 3x + 7 }{ \sqrt{x}} dx\] Using u-sub?

ganeshie8 (ganeshie8):

\(\sqrt{x} = u\)

OpenStudy (anonymous):

Try using \[ \int x^n\;dx = \frac{x^{n+1}}{n+1} \]As well as: \[ \int f(x)+g(x)\;dx = \int f(x)\;dx+\int g(x)\;dx \]

OpenStudy (shamil98):

I don't see how i'm supposed to use the sum rule for this one wio o.O ganeshie i also don't see how i can do this with u = sqrt(x)..

OpenStudy (anonymous):

\[ \frac{ x^2 + 3x + 7 }{ \sqrt{x}} =\frac{x^2}{\sqrt x}+\frac{3x}{\sqrt x}+\frac{7}{\sqrt x} \]

OpenStudy (shamil98):

oh. well this is going to a lot of work..

ganeshie8 (ganeshie8):

u = sqrt(x) du = 1/sqrt(x) dx

ganeshie8 (ganeshie8):

u^2 = x

ganeshie8 (ganeshie8):

both methods are same work...

ganeshie8 (ganeshie8):

* du = 1/2sqrt(x) dx

OpenStudy (anonymous):

\[ \frac{a^n}{\sqrt x} = ax^n(x^{-1/2})=ax^{n-1/2} \]

OpenStudy (shamil98):

I'm supposed to use u-sub for this one though. so u = sqrt(x) du = 1/sqrt(x) dx Could you walk me through this with u-sub?

OpenStudy (anonymous):

But I don't see any U sub that makes this easier than it already is.

ganeshie8 (ganeshie8):

*du = 1/2sqrt(x) dx

ganeshie8 (ganeshie8):

\(\large u = \sqrt{x} \) \(\implies \) \(\large du = \frac{1}{2\sqrt{x}}dx \) \(\large x = u^2\)

ganeshie8 (ganeshie8):

substitute them in the integral

ganeshie8 (ganeshie8):

^^as wio said, both methods involve same effort

ganeshie8 (ganeshie8):

\(\int\limits_{}^{} \frac{ x^2 + 3x + 7 }{ \sqrt{x}} dx \) changes to \(2\int\limits_{}^{} u^4 + 3u^2 + 7~ du \)

OpenStudy (isaiah.feynman):